A told B " I am thrice as old as you were when I was as old as you are"... If the sum of their present ages is 25... Find their present ages......
Answer by macston(5194)

A=A's age; B=B's age
A=3(B-(A-B))
A=3(2B-A)
A=6B-3A
4A=6B
A=6/4B
.
A+B=25
A=25-B
.
25-B=6/4B
25=10/4B
10=B ANSWER 1: B is 10 years old.
.
A+B=25yrs
A=25yrs-10yrs
A=15yrs ANSWER 2: A is 15 years old.